OpenStudy (anonymous):

1help step by step please on how to do this 5^x=4^x+1

OpenStudy (jamesj):

Well, clearly x = 1 is a solution, yes?

OpenStudy (zehanz):

Did you mean: 5^x=4^(x+1)

OpenStudy (anonymous):

yes

OpenStudy (zehanz):

The you could also write :\[5^x=4^{x+1} \Leftrightarrow 5^x=4^x \cdot 4^1 \]

OpenStudy (jamesj):

yes to \[ 5^x = 4^{x+1} \] ?

OpenStudy (anonymous):

yes @JamesJ

OpenStudy (zehanz):

Now you can divide by 4^x

OpenStudy (zehanz):

And use \[\frac{ a^x }{ b^x }=\left( \frac{ a }{ b } \right)^x\]

OpenStudy (anonymous):